### v2

@@ Abstract
We numerically examine dynamo action generated by a flow of an electrically conducting fluid in a precessing cylindrical cavity. We compare a simplified kinematic approach based on the solution of the magnetic induction equation with a prescribed velocity field with the results from a self-consistent three-dimensional simulation of the complete set of magnetohydrodynamic equations. The different kinematic models show a strong influence of the electromagnetic properties of outer layers, which manifests itself in the existence of two different branches with dynamo action. In all cases, we observe a minimum for the onset of dynamo action in a transitional regime, within which the hydrodynamic flow undergoes a change from a large-scale to a more small-scale, turbulent behaviour. However, significant differences in the absolute values for the critical magnetic Reynolds number occur, and they can be explained by the low rotation rate in the MHD simulations achievable for technical reasons. In these models the character of the dynamo solution is small-scale and the magnetic energy remains significantly smaller than the kinetic energy of the flow. In irregular intervals, we observe dynamo bursts with a local concentration of the magnetic field, resulting in a global increase of the magnetic energy by a factor of 3 to 5. However, diffusion of the local patches caused by strong local shear is too rapid, causing these features to exist for only a short period (i.e., shorter than a rotation period) so that their dynamical impact on the dynamo remains small. As the magnetic field is small-scale and weak, the nonlinear feedback on the flow through the Lorentz force remains small and arises essentially in terms of a slight damping of the fast timescales, whereas there is no noticeable change in flow amplitude compared to the hydrodynamic case.
@@ I Introduction
The experimental investigation of magnetic field generation through a flow of electrically conductive fluid is of great interest, as such a dynamo process is crucial for many astrophysical bodies and allows conclusions to be drawn about their evolution and internal structure. Following the successful first-generation dynamo experiments in Karlsruhe [1] and Riga [2], which confirmed the fundamental principle, and subsequent dynamo experiments in Cadarache [3], characterized by a wide variety of different dynamic behaviours, a new experiment is currently being set up at HZDR [4, 5]. In this experiment, a flow of liquid sodium in a precessing cylinder is intended to generate a magnetic field. A mechanical drive such as precession is indeed capable of generating strong flows, allowing for the large flow amplitudes required for the occurrence of the dynamo effect. This contrasts with potentially more straightforward mechanisms, such as convection, which at typical laboratory scales may not achieve sufficiently vigorous flows due to lower driving efficiency. In addition, precession as a flow drive makes perfect sense in a dynamo experiment because driving on the largest possible length scale – the size of the container – opens up significantly more degrees of freedom for establishing a (turbulent) flow through cascade processes than in the case of the previously highly optimized experiments[6].
Precession is a repeatedly proposed driving mechanism for natural dynamos, whether for Earth [7] or the Moon [8]. Paleomagnetic records suggest that the geodynamo was active even before the formation of the Earth’s solid inner core, which implies that thermochemical convection, driven by the cooling and crystallization of the core, was not always the primary mechanism. Before the onset of thermochemical convection, alternative mechanisms must have driven the early geodynamo. Among the most plausible candidates are precession-driven and tide-driven flows, which could have provided the necessary energy for maintaining the magnetic field in the Earth’s early history. This follows from recent experiments and numerical computation yield rather large values for the heat conductivity under conditions typical for the Earth’s liquid core, which imposes strict energetic constraints for the convective state, particularly for the early geodynamo (before the formation of the solid inner Earth’s core, see e.g. Olson [9], Landeau et al. [10]). While it is relatively clear that the laminar flow directly driven by precession is energetically insignificant [11], the original idea of Malkus was based on a turbulent base state [7], which was underpinned by estimates of an upper limit to (viscous) dissipation by Kerswell [12]. Indeed, precession-driven fluid flows exhibit several instabilities, allowing both forward and backward cascades of energy transfer [13] that end up in a flow state with small-scale turbulence superimposed by large-scale flow contributions essentially in terms of a circulation opposite to the original rotation of the container [14, 15].
Regarding the Lunar dynamo, which around 4 billion years ago generated a magnetic field of comparable strength to that of today’s Earth, precession is indeed the most promising candidate to explain the observations including the disappearance of the dynamo about 3 billion years ago, see e.g. Stys and Dumberry [16].
Precession-driven dynamos have been the subject of numerous numerical studies, and there are now successful dynamo simulations in all conceivable geometries, including spheres [17], spheroids [18], cylinders [19], or cubes [20]. A preliminary experiment with sodium in a precessing cylinder was conducted in the early 1970s [21]. This experiment demonstrated the amplification of an externally applied magnetic field by a factor of three, however, the setup at that time was not large enough to achieve the magnetic Reynolds numbers required for a dynamo.
Preliminary investigations using a kinematic model based on the geometry of the planned experiment and the large-scale flow from hydrodynamic simulations showed that the dynamo effect works best in the narrow transition region that separates the subcritical from the supercritical state [4, 22]. In the present study, we deepen the investigation of the dynamo effect in the transitional regime and furthermore, we investigate the influence of the magnetic boundary conditions and turn to full self-consistent models to study the influence of the nonlinear terms and the Lorentz force feedback on the flow.
@@
The present study is organized as follows. In section II.1, we introduce the equations that describe an electrically conducting fluid flow in a precessing cylinder and the related magnetic induction. We further summarize the essential response of the fluid to the forcing imposed by the precession of the container and present the evolution when increasing this forcing. The subsequent section II.3 is dedicated to the results obtained from kinematic dynamo models, where only the induction equation is solved numerically assuming a prescribed, time-independent flow. In this section, we detail the impact of the electrical conductivity of the container. Section III focuses on the corresponding results of the full set of magnetohydrodynamic (MHD) equations and reveals differences and similarities with the kinematic dynamo. We end our study with the conclusions in section IV, where we try to adopt our results for the forthcoming dynamo experiment at HZDR.
II Kinematic approach
II.1 Setup & Equations
@@
Initially, we focus on the kinematic problem, where only the magnetic induction equation is solved and a time-independent velocity field is prescribed, which is supposed to represent the mean (time-averaged) velocity in the experiment. The main finding from this model is a value for the critical magnetic Reynolds number, which is related to the minimum amplitude of the velocity field required for a dynamo to occur. The kinematic approach does neither consider the back-reaction of the magnetic field on the fluid flow via the Lorentz force so that we have no possibility for any saturation, nor the impact of small-scale fluctuations that could hamper (turbulent diffusion) or support (-effect) the dynamo process.
@@
In Eq. (1) represents the velocity field and is the magnetic Reynolds number defined with the angular velocity of the cylinder wit radius and the magnetic diffusivity , which is related to the electrical conductivity by (see also Figure 1a).
In all cases, the inner region of the simulation domain consists of a cylindrical container with radius and height (in dimensionless unities) which is filled with an electrically conducting fluid with a magnetic diffusivity . The fluid flow is prescribed by the data taken from hydrodynamic simulations presented in Pizzi, Giesecke, and Stefani [23], Pizzi et al. [13], Giesecke et al. [15]. In these studies, it is found that the flow in a cylinder emerging at a large nutation angle can be represented by the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow. A more detailed prescription of the structure of the flow field is given in the following subsection.
The flow fields used in this study are derived from hydrodynamic simulations with the spectral element code SEMTEX[24, 25]. A detailed description of the flow field as a function of the precession ratio defined by can be found in [23, 13, 14, 15]. In the following, we briefly summarize the main properties that are relevant for our present study. The flow in a cylinder precessing at a large nutation angle can be characterized by three regimes. In all cases, it is possible to represent the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow. The structure of the time-averaged flow fields as applied in the kinematic model is shown in Figure 2 for three precession ratios that are characteristic for the three flow regimes.
In the subcritical regime, the flow is essentially laminar, and the geometry can be described by the superposition of the rotational motion due to the rotating cylinder and the directly driven flow (essentially an mode proportional to , see left column in Figure 2). Overlaying this is the zonal flow maintained by the cylinder’s rotation (geostrophic and axisymmetric). Without precession, this would be a pure solid-body rotation with with but due to the influence of precession the radial profile of the zonal flow undergoes a significant modification which leads to a significant deceleration of the fluid rotation in the bulk (see bottom row in Figure 2). For larger forcing the flow transients into a supercritical region, in which the large-scale flow in the inner region has almost completely stopped so that the fluid essentially performs small-scale turbulent fluctuations and large-scale components are then only found near the walls of the container (see right column in Figure 2). These two regimes are linked by a transition zone, which is characterized by an intermittent behaviour of the large-scale flow with elements from both regimes. In this transitional region, there exists an additional large-scale contribution in the form of a double roll pattern as shown in the central column of Figure 2 (third row).
II.2 Code
II.2.1 Algorithm
@@
Assuming a prescribed time-independent flow field, Eq. (1) specifies a linear problem and the solution follows with the complex eigenvalue consisting of the growth rate (real part ) and the frequency (imaginary part ). In order to solve Eq. (1) numerically, we use a finite volume approach in cylindrical coordinates. The scheme applies the constraint transport method to guarantee divergence-free solutions and is described in more detail in Giesecke, Stefani, and Gerbeth [26]. Extensive tests of the algorithm, in particular the impact of locally varying material properties, such as electrical conductivity and comparison with an alternative approach based on Spectral/Finite Elements (the SFEMaNS code) can be found in Giesecke et al. [27].
II.2.2 Boundary conditions
We apply pseudo-vacuum conditions for the magnetic field, which read (, where is the unit normal vector at the boundary so that the tangential components of the magnetic field an the outer boundary vanish. It is well known that pseudo-vacuum conditions in kinematic models usually underestimate the critical magnetic Reynolds number that must be exceeded for the dynamo to start. The reduction depends on the geometry, i.e. essentially on the aspect ratio, and amounted, for example, to up to 30% in the kinematic models of the VKS dynamo compared to insulator boundary conditions [27]. Nevertheless, we must restrict ourselves to the application of pseudo-vacuum conditions at the outer boundary of our system, because the strong shear near the outer boundaries enforces a high spatial resolution, which rules out the application of the Boundary-Element-Method (BEM, see e.g. Giesecke, Stefani, and Gerbeth [26]) for the consideration of insulating boundary conditions, as this method requires a large fully populated matrix to be kept in the memory. The size of this matrix scales with , where represents the number of grid points on the cylinder surface so that even current computers with a memory capacity of more than 1 TB are not sufficient to calculate the corresponding models. Instead, we estimate the impact of the magnetic boundary conditions employing various setups with one or two outer layers with thickness and as sketched in Figure 1. The underlying idea is that a sufficiently thick boundary layer with poor electrical conductivity (large magnetic diffusivity) provides a good shielding of the inner volume filled with the electrically conductive fluid so that the actual nature of the electrical boundary conditions sufficiently far away from the induction region is no longer important.
@@ II.2.2 Boundary conditions
We examine three different geometric configurations. Initially we revise the results for a setup that considers only the fluid interior (i.e. ). The second case examines the impact of a container wall with different values of the magnetic diffusivity (i.e. ) and two different values for the thickness of the wall layer. In the thin layer model we use with a resolution of grid cells and in the thick layer model we use with a resolution of grid cells, whereby the magnetic diffusivity in the outer layer remains fixed at the maximum value of . Finally, in the fourth model we again apply a thin cylinder wall with and add a second outer (non-rotating) layer with and , which is supposed to resemble a non-conducting exteriour.
II.3 Results of the kinematic models
II.3.1 General overview
@@
Previous kinematic models demonstrated that dynamo action is most effective when utilizing flow from the transitional region between the subcritical and supercritical states. A detailed investigation of which components of this realization are of decisive importance for the dynamo effect was carried out in Giesecke et al. [22] and showed that shear adjacent to the sidewall plays an important role as well as the presence of the double roll mode (DRM) which seems to be responsible for the considerable drop in the critical magnetic Reynolds number.
II.3.2 No outer layers
@@
The typical behavior of the temporal evolution of the magnetic energy is shown in Figure 3a. Here we focus on two paradigmatic cases with the flow obtained at scaled to a magnetic Reynolds number of (red curve) and (blue curve), which is slightly below and above the dynamo threshold. A striking feature is the oscillating behavior of the energy, which is superimposed on the exponentially growing part. The oscillating character of the solution can be clearly seen in the animation, which can be found in the supplementary material to this study [28].
A systematic study of the impact of the velocity fields obtained at different the critical magnetic Reynolds number required for the onset of dynamo action which is shown in Figure 3b. We do not find any dynamos (at least up to ) for the flow fields in the laminar regime (i.e. below ). The lowest possible magnetic Reynolds number, is , which is obtained for the flow field from hydrodynamic simulations at . It is remarkable, that we find a significant smaller value for in the range than above, which coincides exactly with the transition regime of the velocity field from the subcritical to the supercritical state. We therefore assume that the axially symmetric double roll occurring exclusively in this parameter range plays a significant role for the onset of dynamo action at comparatively small magnetic Reynolds numbers.
II.3.3 Thin outer layer
In the following we focus on the velocity field that provides the minimum critical magnetic Reynolds number in our first models so that all results discussed for the kinematic approach in the reminder of this study are calculated from the velocity field obtained at (central column in Figure 2). We continue with the setup that considers an additional outer layer with thickness with increased magnetic diffusivity that emulates the container wall.

### v3

@@ Abstract
We numerically examine dynamo action generated by a flow of an electrically conducting fluid in a precessing cylindrical cavity. We compare a simplified kinematic approach based on the solution of the magnetic induction equation with a prescribed velocity field with the results from a self-consistent three-dimensional simulation of the complete set of magnetohydrodynamic equations.
In all cases, we observe a minimum for the onset of dynamo action in a transitional regime, within which the hydrodynamic flow undergoes a change from a large-scale to a more small-scale, turbulent behaviour. However, significant differences in the absolute values for the critical magnetic Reynolds number occur depending on the physical properties of the external layers surrounding the flow active domain. The strong influence of the electromagnetic properties of outer layers with the large variation of the critical magnetic Reynolds number can be related to the existence of two different branches with dynamo action.
In contrast to the kinematic models, the nonlinear MHD simulations reveal a small scale dynamo solution with the magnetic energy remaining significantly smaller than the kinetic energy of the flow. In irregular intervals, we observe dynamo bursts with a local concentration of the magnetic field, resulting in a global increase of the magnetic energy by a factor of 3 to 5. However, diffusion of the local patches caused by strong local shear is too rapid, causing these features to exist for only a short period so that their dynamical impact on the dynamo remains small. As the magnetic field is small-scale and weak, the nonlinear feedback on the flow through the Lorentz force remains small and arises essentially in terms of a slight damping of the fast timescales, whereas there is no noticeable change in flow amplitude compared to the hydrodynamic case.
A connection with the kinematic models can be derived by looking at the time-averaged field of the MHD dynamo solution. This is comparable to the eigenmode of the inefficient branch of the kinematic models, which explains their large critical magnetic Reynolds number.
@@ I Introduction
The experimental investigation of magnetic field generation through a flow of electrically conductive fluid is of great interest, as such a magnetohydrodynamic dynamo process is crucial for many astrophysical bodies and allows conclusions about their evolution and internal structure. Following the successful first-generation dynamo experiments in Karlsruhe [1] and Riga [2], which confirmed the fundamental principle, subsequent dynamo experiments in Cadarache (Monchaux et al. [3]) revealed a wide variety of different dynamic behaviours. Currently, a new experiment is being set up at HZDR (Stefani et al. [4]). In the DRESDYN experiment, a flow of liquid sodium in a precessing cylinder is intended to generate a magnetic field. Precession is a repeatedly proposed driving mechanism for natural dynamos, whether for the Moon [5] or for the Earth [6]. Regarding the Lunar dynamo, which around 4 billion years ago generated a magnetic field of comparable strength to that of today’s Earth, precession is indeed the most promising candidate to explain the observations including the disappearance of the dynamo about 3 billion years ago (see e.g. Stys and Dumberry [7]). One of the main reasons to consider precession driven dynamo action for the geodynamo are paleomagnetic records, which suggest that the geodynamo was active even before the formation of the Earth’s solid inner core. This, in turn, implies that thermochemical convection, driven by the cooling and crystallization of the core, was not always the primary mechanism and before the onset of thermochemical convection, alternative mechanisms must have driven the early geodynamo. Support for this hypothesis comes from recent experiments and numerical computations that yield rather large values for the heat conductivity under conditions typical for the Earth’s liquid core, which imposes strict energetic constraints for the convective state, particularly for the early geodynamo before the formation of the solid inner Earth’s core (see e.g. Olson [8], Landeau et al. [9]). While it is relatively clear that the laminar flow directly driven by precession is energetically insignificant (Rochester et al. [10]), the original idea of Malkus was based on a turbulent base state[6], which was later underpinned by estimates of an upper limit to (viscous) dissipation by Kerswell [11]. Indeed, precession-driven fluid flows exhibit several instabilities, allowing both forward and backward cascades of energy transfer (Pizzi et al. [12]) that end up in a flow state with small-scale turbulence superimposed by large-scale flow contributions essentially in terms of a circulation opposite to the original rotation of the container (Wilbert, Giesecke, and Grauer [13], Giesecke et al. [14]). This behaviour has also been observed in the laboratory[15] and so far various experiments have shown that precession is capable of generating strong flows, principally allowing for the large flow amplitudes required for the occurrence of the dynamo effect[16, 17, 18, 19, 20, 21, 22, 23, 24]. This contrasts with potentially more straightforward mechanisms, such as convection, which at typical laboratory scales may not achieve sufficiently vigorous flows due to lower driving efficiency [25, 26, 27]. Precession-driven dynamos also have been the subject of numerous numerical studies, and there are now successful dynamo simulations in all conceivable geometries, including spheres [28], spheroids [29], or cubes [30]. Nore et al. [31] and Cappanera et al. [32] investigated dynamo solutions in a cylindrical geometry comparable to the DRESDYN experiment, but with small Reynolds numbers () and a rather large precession ratio of . The induced magnetic fields showed a complex structure, which correspond most closely to a quadrupolar geometry. Preliminary investigations with particular regard to the DRESDYN dynamo experiment utilized a kinematic model based on the geometry of the planned experiment and the large-scale flow from hydrodynamic simulations (Giesecke et al. [33, 34]). These models showed that the dynamo effect works best in the narrow transition region that separates the subcritical from the supercritical state (Pizzi et al. [35], Kumar et al. [36]). In the present study, we deepen the investigation of the dynamo effect in the transitional regime, investigate the influence of outer non-fluid layers, and compare kinematic models with a self-consistent magnetohydrodynamic (MHD) model to study the influence of the nonlinear terms and the Lorentz force feedback on the flow.
@@
The present study is organized as follows. In section II, we introduce the equations that describe an electrically conducting fluid flow in a precessing cylinder and the related magnetic induction. We further summarize the essential response of the fluid to the forcing imposed by the precession of the container and present the evolution when the forcing is increased. The subsequent section III is dedicated to the results obtained from kinematic dynamo models, where only the induction equation is solved numerically assuming a prescribed, time-independent flow. In this section, we detail the impact of the electrical conductivity of the container. Section IV focuses on the corresponding results obtained from the full set of magnetohydrodynamic (MHD) equations and reveals differences and similarities with the kinematic dynamo. We end our study with the conclusions in section V, where we try to adopt our results for the forthcoming dynamo experiment at HZDR.
II Setup of the kinematic dynamo models
II.1 Equations and numerical methods
@@
Initially, we focus on the kinematic problem, where only the magnetic induction equation is solved, and a time-independent velocity is prescribed, which is supposed to represent the mean (time-averaged) velocity in the experiment. Here we compare different configurations without and with outer layers of a conductive material surrounding the actual medium on which the force is exerted by the precession. The main finding from these models is a value for the critical magnetic Reynolds number, which is related to the minimum amplitude of the velocity field required for a dynamo to occur. The kinematic approach does neither consider the back-reaction of the magnetic field on the fluid flow via the Lorentz force nor the impact of small-scale fluctuations that could obstruct (turbulent diffusion) or support (-effect) the dynamo process.
@@
In Eq. (1) represents the (incompressible) velocity field and is the magnetic Reynolds number defined with the angular velocity of the cylinder, the radius and the magnetic diffusivity (see also Figure 1a).
@@
Assuming a prescribed time-independent flow field, Eq. (1) specifies a linear problem and the solution follows with the complex eigenvalue consisting of the growth rate (real part ) and the frequency (imaginary part ). In order to solve Eq. (1) numerically, we use a finite volume approach in cylindrical coordinates. The scheme applies the constraint transport method to guarantee divergence-free solutions and is described in more detail in Giesecke, Stefani, and Gerbeth [37]. Extensive tests of the algorithm, in particular the impact of locally varying material properties, such as electrical conductivity and comparison with an alternative approach based on Spectral/Finite Elements (the SFEMaNS code[38]) can be found in Giesecke et al. [39].
II.2 Boundary conditions and nested outer layers
@@
We estimate the impact of the magnetic boundary conditions employing various setups with one or two outer layers with thickness and as sketched in Figure 1b. The underlying idea is that a sufficiently thick outer layer with poor electrical conductivity (large magnetic diffusivity) reduces the impact of the magnetic boundary conditions on the inner volume filled with the electrically conductive fluid. We consider several models that differ in the presence of up to two outer layers with thicknesses and that enclose the internal region of the container filled with an electrically conducting fluid (Figure 1b). In all cases, the inner region of the simulation domain consists of a cylindrical container with radius and height (in dimensionless units) which is filled with an electrically conducting fluid with a magnetic diffusivity . We examine three different geometric configurations. Initially we revise the results for a setup that considers only the fluid interior (i.e. ). The second case examines the impact of a container wall with different values of the magnetic diffusivity (i.e. ) and two different values for the thickness of the wall layer. In the thin layer model we use with a resolution of grid cells and in the thick layer model we use with a resolution of grid cells, whereby the magnetic diffusivity in the wall layer remains fixed at the maximum value of . As we tackle the problem in the precession frame of reference (with the observer co-rotating with the precession while looking at the spinning container), the wall layer follows a rotational motion with and .
Finally, in the fourth model we again apply a thin wall layer with and surrounded by a second, outer layer with and , which is supposed to resemble a non-conducting exterior (see Rädler et al. [40] and present discussion on page 14). This layer does not co-rotate with the container so that the velocity is set to zero within this layer. At the outer boundaries of the computational domain we apply pseudo-vacuum conditions for the magnetic field, which read (, where is the unit normal vector at the boundary so that the tangential components of the magnetic field an the outer boundary vanish. It is well known that pseudo-vacuum conditions in kinematic models usually underestimate the critical magnetic Reynolds number that must be exceeded for the dynamo to start. The reduction depends on the geometry, i.e. essentially on the aspect ratio, and amounted, for example, to up to 30% in the kinematic models of the VKS dynamo compared to realistic insulator boundary conditions (Giesecke et al. [39]).
II.3 Velocity field
The fluid flow is prescribed using the data taken from hydrodynamic simulations presented in Pizzi, Giesecke, and Stefani [41], Pizzi et al. [12], Wilbert, Giesecke, and Grauer [13] and Giesecke et al. [14]. In these studies, it was found that the flow in a cylinder emerging at a large nutation angle can be represented by the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow. In dependence of the precession ratio defined by the flow basically can be characterized by three regimes, which we simply name the subcritical and the supercritical state, and the transition section between both states. A compact representation of the three regimes, which illustrates how the transition between the subcritical state and the supercritical state takes place, is provided in Figure 2, which shows the ratio of the energy of the turbulent flow component in the form of the time- and volume-averaged root-mean-square velocity to the energy of the actual axisymmetric geostrophic component as a function of . The plot illustrates the abrupt shift from the rotation-dominated regime (below ) to the turbulence-dominated regime (above ).
However, although the fluctuations of the flow with respect to the fluid rotation increase strongly with , the dominant energetic contributions of the precession-induced flow can always be captured by a few large-scale inertial modes, since the significant jump in Figure 2 is determined by both, the increase in turbulence intensity, and the deceleration of the fluid rotation in the core of the cylinder (see Giesecke et al. [14]).
The typical structure of the time-averaged flow fields is shown in Figure 3 for three precession ratios that are characteristic for the three flow regimes.
In the subcritical state (left column in Figure 3), the flow is essentially laminar, and the geometry can be described by the directly driven flow (first and second row, essentially an mode proportional to ) and a zonal flow maintained by the cylinder’s rotation (geostrophic and axisymmetric, third row). Without precession, this would be a pure solid-body rotation with but due to the influence of precession the radial profile of the zonal flow is significantly modified (i.e. deceleration of the fluid rotation in the bulk, see bottom row in Figure 3). For larger forcing the flow transients into a supercritical region, in which the large-scale flow in the inner region almost vanishes so that the fluid essentially performs small-scale turbulent fluctuations and large-scale components are only found near the walls of the container (see right column in Figure 3). These two regimes are linked by a transition zone, which is characterized by an intermittent behavior of the large-scale flow with elements from both regimes. In this transitional region, there exists an additional large-scale contribution in the form of a double roll pattern as shown in the third row of the central column of Figure 3.
@@ II.3 Velocity field
This double roll is rather similar to the (time-averaged) poloidal velocity field in the VKS dynamo111In that dynamo experiment the flow was driven by two oppositely rotating disks, and it is well known that such a flow structure generates a dynamo with a relatively low critical magnetic Reynolds number[42] (when also the associated azimuthal flow is taken into account). Indeed, previous kinematic models using a time-averaged flow driven by precession demonstrated that dynamo action is most effective when utilizing the flow obtained in the transitional region between the subcritical and supercritical states [33, 34]. A detailed investigation of which components (in terms of inertial modes) are of decisive importance for the dynamo effect was carried out in Giesecke et al. [34] and showed that shear adjacent to the side walls plays an important role as well as the presence of the double-roll mode which seems to be responsible for the considerable drop in the critical magnetic Reynolds number.
III Kinematic growth rates and critical magnetic Reynolds number
III.1 General overview
So far kinematic models for the DRESDYN dynamo experiment mostly applied pseudo-vacuum conditions for the magnetic field and only few models have been published where an outer wall with finite thickness and with different magnetic diffusivity has been considered [34]. These models indicate that a container wall made of stainless steel, i.e. with a diffusivity that is larger by one order of magnitude when compared to the liquid sodium in the interior, yields growth rates for the magnetic energy similar to the models with pseudo-vacuum conditions. In the following, we will review and deepen these results.
III.2 No layer
@@
The typical behavior of the temporal evolution of the magnetic energy is shown in Figure 4a. Here we focus on two paradigmatic cases with the flow obtained at scaled to a magnetic Reynolds number of (red curve, decaying solution) and (blue curve, growing solution), which are slightly below and above the dynamo threshold. A striking feature is the oscillating component of the energy, which is superimposed on the exponentially growing part. For a flow field with a slightly increased , the behavior looks similar in principle, but then a higher magnetic Reynolds number is required for the onset of a dynamo, and the frequency of the oscillating component increases as shown by the orange and green curves in Figure 4a, which result from the flow field obtained at .
The distribution of the magnetic energy looks similar in both cases (see Figure 4b). It is worth mentioning in relation to results including outer boundary layers (see below) that a significant part of the magnetic field energy is observed in the bulk of the cylinder. The temporal dynamics due to the oscillating field component are shown in an animation, which can be found in the supplementary material to this study [43].
III.3 Wall layer
We continue with the setup that considers an additional wall layer with thickness . Within this layer the magnetic diffusivity is increased in order to emulate the properties of a container wall made of stainless steel. With such an outer layer, we essentially find two new features: (1) the onset for dynamo action is shifted to larger magnetic Reynolds numbers
and (2) we can distinguish two different types of dynamo action: if the wall diffusivity is close to that of the fluid (i.e. ), a dynamo only occurs at very large (). This dynamo has no oscillatory components (see orange and green curve in Figure 5a) and the associated field energy is concentrated in a narrow region near the end caps (top plot with iso-surfaces in Figure 5b). If the wall layer has an (see red and blue curves in Figure 5a), we obtain an oscillating solution that is similar to the solution in the previous case with pseudo-vacuum boundary conditions. In that case the magnetic field energy is again widely distributed throughout the entire volume with the characteristic pattern shown in the bottom plot of Figure 5b.